# Dynamic Pricing with Adversarially-Censored Demands

### v1

@@ 5 Regret Analysis
In this section, we analyze the cumulative regret of our algorithm and show a $\tilde{O}(\sqrt{T})$ regret guarantee with high probability. We leave all of the proof details to Appendix A, and here we only display key lemmas and proof sketches.
We firstly propose our main theorem.
Theorem 5.1 (Regret).
Let $\tau=\frac{1}{\sqrt{T}}$ in Algorithm 1. For any adversarial $\{\gamma_{t}\}_{t=1}^{T}$ input sequence, C20CB suffers at most $\tilde{O}\left(\sqrt{T}\cdot\log\frac{T}{\delta}\right)$ regret, with probability $\Pr\geq 1-\delta$.
Proof.
In order to prove Theorem 5.1, we have to show the following three components:
@@
In the following, we present each corresponding lemma regarding to the roadmap listed above.
Lemma 5.2 (revenue function $r_{t}(p)$).
For the expected revenue function $r_{t}(p)$ defined in eq. 2, the following properties hold:
@@
    $r_{t}(p)$ is unimodal: There exists a unique $p_{t}^{*}\in[0,\frac{a}{b}]$ such that $r^{\prime}_{t}(p_{t}^{*})=0$, and $r_{t}(p)$ monotonically increase in $[0,p_{t}^{*}]$ and decrease in $[p_{t}^{*},\frac{a}{b}]$. Notice that $\frac{a}{b}>p_{\max}$ according to Assumption 3.7.
@@
Please kindly check Section A.1 as a detailed proof of Lemma 5.2. The properties of $r_{t}(p)$ and $r^{\prime}_{t}(p)$ enable us upper bound the cost of estimation error and decision bias. In the following, we propose a lemma that serves as a milestone of estimation error upper bounds.
Lemma 5.3 (Estimation error of $a$ and $b$).
With probability $\Pr\geq 1-2\eta\delta$, we have
$|\hat{b}-b|\leq$ $C_{b}\cdot\frac{1}{\sqrt{\tau}},$ (9)
$|\hat{a}-a|\leq$ $C_{a}\cdot\frac{1}{\sqrt{\tau}},$
@@
The key observation to prove this lemma lies in the expectation of each $e_{i,t}$ that indicates whether the demand exceeds certain level under uniformly distributed prices. With the help of a method-of-moment estimate as introduced in Section 4.2, we get rid of the influence of noise distribution and achieves an unbiased estimator of $\frac{1}{b}$ (and therefore $\hat{b}$). With $\hat{b}$ serving as a plug-in estimator, we later get $\hat{a}$. By applying Hoeffding’s Inequalities, we obtain those error bounds. We defer the detailed proof of Lemma 5.3 to Section A.2. With the help of Lemma 5.3, we may upper bound the estimation error of $r^{\prime}_{t}(p)$ at discrete prices. The error bound is displayed as the following lemma.
Lemma 5.4 (Estimation error of $r^{\prime}_{t}(p_{k,t})$).
There exists constants $C_{N}>0,C_{\tau}>0$ such that for any $t\in[T],k\in\{-M,-M+1,\ldots,M-1,M\}$, its holds with probability $\Pr\geq 1-6\eta\delta$
@@
We kindly refer the readers to Section A.3 as a rigorous proof of Lemma 5.4. Given this lemma, the derivatives of each discrete price $p_{k,t}$ is truthfully reflected by their corresponding error bound. Therefore, we intuitively see that the closest-to-zero confidence bound represents the closest-to-$p_{t}^{*}$ discrete price. We formulate this intuition as the following lemma:
Lemma 5.5 (Closest-to-zero confidence to performance).
@@
    When $\exists k\in\{-M,-M+1,\ldots,M-1,M\}$ such that the Number $k$ confidence bound satisfies $\hat{r}_{k,t}-\Delta_{k}(t)\leq 0\leq\hat{r}_{k,t}+\Delta_{k}(t)]$, and also $N_{k}(t)>N_{0}$, then we have $p_{k,t}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}]$. Furthermore, there exists constant $C_{in}$ such that $r_{t}(p_{t}^{*})-r_{t}(p_{k,t})\leq C_{in}(\frac{1}{N_{k}(t)}+\frac{1}{\tau})$.
@@
    When there exists no confidence bound that contains 00, i.e. either $\hat{r}_{k,t}-\Delta_{k}(t)>0$ or $\hat{r}_{k,t}+\Delta_{k}(t)<0,\forall k\in\{-M,-M+1,\ldots,M-1,M\}$ (happens at least for one $k$), and also $N_{k}(t)>N_{1}$, then we have
    $\inf_{k}\min\{|\hat{r}_{k,t}-\Delta_{k}(t)|,|\hat{r}_{k,t}+\Delta_{k}(t)|\}\leq\frac{L_{r}(C_{a}+C_{b}\cdot p_{\max})}{2b_{\min}}\cdot\frac{1}{\sqrt{\tau}},$
    and also $p_{k,t}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}]$ Furthermore, there exists constant $C_{out}$ such that $r_{t}(p_{t}^{*})-r_{t}(p_{k,t})\leq C_{out}(\frac{1}{N_{k}(t)}+\frac{1}{\tau})$.
The intuition to prove Lemma 5.5 is twofold:
@@
    When an error bar contains 00, the true derivative of the corresponding price is close to 00 within the distance of its error bound. By applying Lemma 5.2 Property (7), we may upper bound the performance loss with the square of its derivatives, which is further upper bounded by the square of error bound.
@@
    When no error bar contains 00, there exists an adjacent pair of prices whose error bars are separated by $y=0$. On the one hand, their derivatives difference is upper bounded due to the Lipschitzness of $r^{\prime}_{t}(p)$. On the other hand, the same derivatives difference is lower bounded by the closest-to-zero confidence bound. Therefore, the gap between $y=0$ and the closest-to-zero confidence bound should be very small, and we still have a comparably small $|r^{\prime}_{t}(p_{t})|$ if $p_{t}$ possesses that confidence bound. As a consequence, we have similar upper bound on the performance loss comparing with Case (1), up to constant coefficients.
The detailed proof of Lemma 5.5 is presented in Section A.4. Finally, we have a lemma that upper bounds the regret of proposing $p_{t}=\frac{\hat{a}}{2\hat{b}}$ under specific circumstances.
Lemma 5.6 (Proposing $\frac{\hat{a}}{2\hat{b}}$).
When $\gamma_{t}>\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c$ and when $\hat{r}_{k,t}-\Delta_{k}(t)>0,\forall k=-M,-M+1,\ldots,M-1,M$, we have $p_{t}^{*}=\frac{a}{2b}$ and there exists a constant $C_{non}$ such that
$r_{t}(\frac{a}{2b})-r_{t}(\frac{\hat{a}}{\hat{2}b})\leq C_{non}\frac{1}{\tau}.$ (11)
The intuition of Lemma 5.6 is that $\frac{a}{2b}$ is the optimal price without censoring, and we only need to show that either the optimal price or $\frac{a}{2b}$ is not censored (which are equivalent as the revenue function is unimodal). We defer its proof to Section A.5. This lemma serves as the last puzzle of the proof. With all lemmas above, we may upper bound the cumulative regret as follows:
$Regret$ $=\sum_{t=1}^{T}r_{t}(p_{t}^{*})-r_{t}(p)$ (12)
$\leq\tau\cdot a_{\max}p_{\max}+(2M+1)(1+N_{0}+N_{1})\cdot a_{\max}p_{\max}$
$\quad+\sum_{t=1}^{T}(\max\{C_{in},C_{out}\})(\frac{1}{N_{k}(t)}+\frac{1}{\tau})+\sum_{t=1}^{T}C_{non}\cdot\frac{1}{\tau}$
[…]
$=\tilde{O}(\sqrt{T}+\sum_{k=1}^{2M+1}\sum_{i_{k}=1}^{N_{k}(T)}\frac{1}{i_{k}})$
$=\tilde{O}(\sqrt{T}+T^{\frac{1}{4}}\log T)$
$=\tilde{O}(\sqrt{T}).$
By applying a union bound, we know that Eq. 12 holds with probability
@@
Here the first part comes from Lemma 5.3, and the second part comes from Lemma 5.4 for any $t\in[T]$ and $k\in\{-M,\ldots,M\}$. Let $\eta:=\frac{C_{a}+C_{b}\cdot p_{\max}}{10c\cdot T^{5/4}}$, and we show that Theorem 5.1 holds. ∎
Remark 5.7.
This $\tilde{O}(\sqrt{T})$ regret upper bound is (near) optimal up to $\log{T}$ factors, as it matches the $\Omega(\sqrt{T})$ information-theoretic lower bound proposed by Broder and Rusmevichientong, (2012) for a no-censoring problem setting with linear noisy demand.
From the analysis above, we notice that the threshold of learning the optimal prices in our problem setting is still the estimation of parameters. The assumptions we made in Section 3 scales the efficiency of learning $\hat{b}$ and $\hat{a}$, which we will discuss in Section 6.
@@ 6 Discussions
Here we discuss the limitations, potential extensions and impacts of our work.
Generalization to Unbounded Noises
We assume the noise is bounded in a constant-width range. This assumption streamlines the pure-exploration phase and facilitates the estimation of the parameters $b$ and $a$. While our methods and results can be extended to unbounded $O(\frac{1}{\log T})$-subGaussian noises by simple truncation, challenges remain for handling generic unbounded noises. Moreover, the problem can be more sophisticated with dual-censoring, both from above by inventory–as we have discussed– and from below by 00, especially when considering unbounded noises.
Extensions to Contextual Pricing
In this work, we assume $a$ and $b$ are static, which may not hold in many real scenarios. Example 1.1 serves as a good instance, showcasing significant fluctuations in popularity across different performances. A reasonable extension of our work would be modeling $a$ and $b$ as contextual parameters. Similar modelings have been adopt by Wang et al., (2025) and Ban and Keskin, (2021) in the realm of personalized pricing research.
Extensions to Non-Lipschitz Noise CDF
In this work, we assume the noise CDF as a Lipschitz function as many pricing-related works did (Fan et al.,, 2024; Tullii et al.,, 2024). This assumption enables the local smoothness at $p_{t}^{*}$ and leads to a quadratic loss. However, this prevent us from applying our algorithm to non-Lipschitz settings, which even includes the noise-free setting. In fact, although we believe that a better regret rate exists for the noise-free setting, we have to state that the hardness of problems are completely different with Lipschitz noises versus without it. Although a Lipschitz noise makes the observation “more blur”, it also makes the revenue curve “more smooth”. We would like to present an analog example from the feature-based dynamic pricing problem: When the Gaussian noise $\mathcal{N}(0,\sigma^{2})$ is either negligible (with $\sigma<\frac{1}{T}$, see Cohen et al., (2020)) or super significant (with $\sigma>1$, see Xu and Wang, (2021)), the minimax regret is $O(\log T)$. However, existing works can only achieve $O(\sqrt{T})$ regret when $\sigma\in[\frac{1}{T},1]$. We look forward to future research on our problem setting once getting rid of the Lipschitzness assumption.
Extensions to Non-linear Demand Curve
@@
In this work, we adopt a linear-and-noisy model for the potential demands, which is standard in dynamic pricing literature. Also, we utilize the unimodal property brought by this linear demand model, even after the censoring effect is imposed. If we would like to generalize our methodologies to non-linear demand functions, especially where the unimodality does not hold any longer, we have to carefully distinguish between local optimals and saddle-points that may also cause $r^{\prime}_{t}(p)=0$ for some sub-optimal $p$. We conjecture an $\Omega(T^{\frac{m+1}{2m+1}})$ lower bound in that case, where $m$ is the time of smoothness. However, it is still unclear whether the censoring effect will introduce new local optimals or swipe off existing ones in multimodal settings.
Societal Impacts
@@
Our research primarily addresses a non-contextual pricing model that does not incorporate personal or group-specific data, thereby adhering to conventional fairness standards relating to temporal, group, demand, and utility discrepancies as outlined by Cohen et al., (2022); Chen et al., (2023); Xu et al., (2023). However, the non-stationarity of inventory levels could result in varying fulfillment rate over time, i.e. the proportions of satisfied demands at $\{p_{t}^{*}\}$’s might be different for $t=1,2,\ldots,T$. This raises concern regarding unfairness in fulfillment rate (Spiliotopoulou and Conte,, 2022) particularly on product of significant social and individual importance.

### v2

@@ 5 Regret Analysis
In this section, we analyze the cumulative regret of our algorithm and show a $\tilde{O}(\sqrt{T})$ regret guarantee with high probability. Here we only display key lemmas and proof sketches, and we leave all of the proof details to Appendix˜0.B.
We first state our main theorem.
Theorem 5.1(Regret)
Let $\tau={\sqrt{T}}$ in Algorithm˜1. For any adversarial $\{\gamma_{t}\}_{t=1}^{T}$ input sequence, C20CB suffers at most $\tilde{O}\left(\sqrt{T}\cdot\log\frac{T}{\delta}\right)$ regret, with probability $\Pr\geq 1-\delta$.
Proof
In order to prove Theorem˜5.1, we have to show the following three components:
@@
In the following, we present each corresponding lemma regarding to the roadmap above.
Lemma 1(Revenue Function $r_{t}(p)$)
For the expected revenue function $r_{t}(p)$ defined in eq.˜2, the following properties hold:
@@
    $r_{t}(p)$ is unimodal: There exists a unique $p_{t}^{*}\in[0,\frac{a}{b}]$ such that $r^{\prime}_{t}(p_{t}^{*})=0$, and $r_{t}(p)$ monotonically increase in $0,p_{t}^{*}]$ and decrease in $p_{t}^{*},\frac{a}{b}]$. Notice that $\frac{a}{b}>p_{\max}$ according to Assumption˜4.
@@
Proof(sketch)
Property 1 is from integration by parts. Property 2 is proved by the Lipschitzness of $F(x)$. Property 3 is from the monotonicity of $F(x)$. Property 4 can be proved by two steps: (4.1) The existence of $p_{t}^{*}\in[0,\frac{a}{b}]$ by $r_{t}^{\prime}(0)>0$ and $r_{t}^{\prime}(\frac{a}{b})<0$; (4.2) The uniqueness of $p_{t}^{*}$ by contradiction. Properties 5 and 6 are mainly from the Lipschitzness of $r^{\prime}_{t}(p)$. Property 7 comes from $r^{\prime}_{t}(p_{t}^{*})=0$ and the strong concavity (Property 6). Please kindly check Section˜0.B.1 as a detailed proof of Lemma˜1. ∎
The properties of $r_{t}(p)$ and $r^{\prime}_{t}(p)$ enable us to upper bound the cost of estimation error and decision bias. In the following, we propose a lemma that serves as a milestone of estimation error upper bounds.
Lemma 2(Estimation Error of $a$ and $b$)
For any $\eta>0,\delta>0$, with probability $\Pr\geq 1-2\eta\delta$, we have
$|\hat{b}-b|\leq$ $C_{b}\cdot\frac{1}{\sqrt{\tau}},\quad|\hat{a}-a|\leq C_{a}\cdot\frac{1}{\sqrt{\tau}},$ (10)
@@
Proof(sketch)
The key observation to prove this lemma lies in the expectation of each $e_{i,t}$ that indicates whether the demand exceeds certain level under uniformly distributed prices. According to Law of Total Expectation, for any $\gamma\in[\gamma_{\min},\gamma_{t}]$ at any time $t\leq\tau$ in STAGE 1, we have
$\mathbb{E}[\mathds{1}[D_{t}\geq\gamma]]=$ $\mathbb{E}_{N_{t}}[\mathbb{E}_{p_{t}}[a-bp_{t}+N_{t}\geq\gamma|N_{t}]]$ (11)
$=$ $\mathbb{E}_{N_{t}}[\mathbb{E}_{p_{t}}[p_{t}\leq\frac{a-\gamma+N_{t}}{b}|N_{t}]]$
$=$ $\mathbb{E}_{N_{t}}[\frac{a-\gamma+N_{t}}{b\cdot p_{\max}}]=\frac{a-\gamma}{bp_{\max}}.$
With this property, we may construct method-of-moment estimates from $e_{i,t}$ in STAGE 1, which eliminate the influence of noise distribution and achieve an unbiased estimator of $\frac{1}{b}$ (and therefore $\hat{b}$ asymptotically). With $\hat{b}$ serving as a plug-in estimator, we later achieve $\hat{a}$. We obtain the error bounds by applying Hoeffding’s Inequalities. ∎
We defer the detailed proof of Lemma˜2 to Section˜0.B.2. With the help of Lemma˜2, we may upper bound the estimation error of $r^{\prime}_{t}(p)$ at discrete prices. The error bound is displayed as the following lemma.
Lemma 3(Estimation Error of $r^{\prime}_{t}(p_{k,t})$)
There exists constants $C_{N}>0,C_{\tau}>0$ such that for any $t\in[T],k\in\{-M,-M+1,\ldots,M\}$, with probability $\Pr\geq 1-6\eta\delta$ we have
@@
Proof(sketch)
Denote $N_{k}(t),F_{k}(t)$ and $G_{k}(t)$ as the value of $N_{k},F_{k}$ and $G_{k}$ at the beginning of time $t$. From Algorithm 1, we have $|\hat{r}_{k,t}-r^{\prime}_{t}(p_{k,t})|\leq|G_{k}(t)-(G(c)-G(\gamma_{t}-a+bp_{k,t}))|+p_{k,t}|\hat{b}F_{k}-bF(\gamma_{t}-a+bp_{k,t})|$, and we may separately upper bound each of these two differences.
  1. 1.
    For the term on $G$, we may split the quantity $|G_{k}(t)-(G(c)-G(\gamma_{t}-a+bp_{k,t}))|$ into three terms:
    1. (i)
      $|G_{k}(t)-\mathbb{E}[G_{k}(t)]|$ as concentration, which is bounded by Hoeffding’s inequality.
    2. (ii)
      $|\mathbb{E}[G_{k}(t)]-(G(c)-G(2k\Delta))|$ as the $G_{k}$ deviation within each “bin” of estimation (whose center is $w_{k}=2k\Delta$). By definition of $G_{k}(t)$, we have $\mathbb{E}[G_{k}(t)]=\frac{1}{N_{k}(t)}\sum_{s=1}^{t-1}\mathds{1}[k_{s}==k](G(c)-G(\gamma_{s}-a+bp_{k_{s},s}))$, and therefore we have
      $|\mathbb{E}[G_{k}(t)]-(G(c)-G(2k\Delta))|$ (13)
      $\leq$ $\frac{1}{N_{k}(t)}\sum_{s<t:k_{s}==k}|G(\gamma_{s}-a+bp_{k_{s},s})-G(2k\Delta)|$
      $\leq$ $O(|\hat{a}-a|+|\hat{b}-b|).$
      The last line is by Lipschitzness of $G(x)$.
    3. (iii)
      $|G(2k\Delta)-G(\gamma_{t}-a+bp_{k,t})|$, which is similar to the last step of (ii) above.
  2. 2.
    For the term on $F$, we may split the quantity $|\hat{b}F_{k}-bF(\gamma_{t}-a+bp_{k,t})|$ into four terms:
    1. (i)
      $|\hat{b}F_{k}(t)-bF_{k}(t)|\leq|\hat{b}-b|$ by the nature of $F_{k}(t)\leq 1$, and further bounded by Lemma˜2.
    2. (ii)
      $|F_{k}(t)-\mathbb{E}[F_{k}(t)]|$, (iii) $|\mathbb{E}[F_{k}(t)]-F(2k\Delta)|$, and (iv) $|F(2k\Delta)-F(\gamma_{t}-a+bp_{k,t})|$, are bounded in the same way presented above for $G$ respectively.
    By plugging in the estimation error from Lemma˜2, we prove the present lemma. ∎
Please refer to Section˜0.B.3 as a rigorous proof of Lemma˜3. Given this lemma, the derivatives of each discrete price $p_{k,t}$ is truthfully reflected by their corresponding error bound. Therefore, we intuitively see that the closest-to-zero confidence bound represents the closest-to-$p_{t}^{*}$ discrete price. We formulate this intuition as the following lemma:
Lemma 4(Closest-To-Zero Confidence To Performance)
@@
    When $\exists k\in\{-M,-M+1,\ldots,M\}$ such that the Number $k$ confidence bound satisfies $\hat{r}_{k,t}-\Delta_{k}(t)\leq 0\leq\hat{r}_{k,t}+\Delta_{k}(t)$, and also $N_{k}(t)>N_{0}$, then we have $p_{k,t}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}]$. Furthermore, there exists constant $C_{in}$ such that $r_{t}(p_{t}^{*})-r_{t}(p_{k,t})\leq C_{in}(\frac{1}{N_{k}(t)}+\frac{1}{\tau})$.
@@
    When there exists no confidence bound that contains 0, i.e. either $\hat{r}_{k,t}-\Delta_{k}(t)>0$ or $\hat{r}_{k,t}+\Delta_{k}(t)<0,\forall k\in\{-M,-M+1,\ldots,M-1,M\}$ (happens at least for one $k$), and also $N_{k}(t)>N_{1}$, then we have
    $\inf_{k}\min\{|\hat{r}_{k,t}-\Delta_{k}(t)|,|\hat{r}_{k,t}+\Delta_{k}(t)|\}\leq C_{\kappa}\cdot\frac{1}{\sqrt{\tau}},$ (14)
    (where $C_{\kappa}=\frac{L_{r}(C_{a}+C_{b}\cdot p_{\max})}{2b_{\min}}$) and $p_{k,t}\in[p_{t}^{*}-\epsilon_{t},p_{t}^{*}+\epsilon_{t}]$. Furthermore, there exists constant $C_{out}$ such that $r_{t}(p_{t}^{*})-r_{t}(p_{k,t})\leq C_{out}(\frac{1}{N_{k}(t)}+\frac{1}{\tau})$.
Proof(sketch)
The intuition to prove Lemma˜4 is twofold:
@@
    When an error bar contains 0, the true derivative of the corresponding price is close to 0 within the distance of its error bound. By applying Lemma˜1 Property (7), we may upper bound the performance loss with the square of its derivatives, which is further upper bounded by the square of error bound.
@@
    When no error bar contains 0, there exists an adjacent pair of prices whose error bars are separated by $y=0$. On the one hand, their derivatives difference is upper bounded due to the Lipschitzness of $r^{\prime}_{t}(p)$. On the other hand, the same derivatives difference is lower bounded by the closest-to-zero confidence bound. Therefore, the gap between $y=0$ and the closest-to-zero confidence bound should be very small, and we still have a comparably small $|r^{\prime}_{t}(p_{t})|$ if $p_{t}$ possesses that confidence bound. As a consequence, we have similar upper bound on the performance loss comparing with Case (1), up to constant coefficients. ∎
The detailed proof of Lemma˜4 is presented in Section˜0.B.4. Finally, we have a lemma that upper bounds the regret of proposing $p_{t}=\frac{\hat{a}}{2\hat{b}}$ under special conditions.
Lemma 5(Proposing $\frac{\hat{a}}{2\hat{b}}$)
When $\gamma_{t}>\frac{\hat{a}+C_{a}\cdot\frac{1}{\sqrt{\tau}}}{2}+c$ and when $\hat{r}_{k,t}-\Delta_{k}(t)>0,\forall k=-M,-M+1,\ldots,M$, we have $p_{t}^{*}=\frac{a}{2b}$ and there exists a constant $C_{non}$ such that
$r_{t}(\frac{a}{2b})-r_{t}(\frac{\hat{a}}{2\hat{b}})\leq C_{non}\frac{1}{\tau}.$ (15)
The intuition of Lemma˜5 is that $\frac{a}{2b}$ is the optimal price without censoring, and we only need to show that either the optimal price or $\frac{a}{2b}$ is not censored (which are equivalent as the revenue function is unimodal). We defer its proof to Section˜0.B.5. This lemma serves as the last puzzle of the proof. With all lemmas above, we upper bound the overall regret as follows:
$Regret=\sum_{t=1}^{T}r_{t}(p_{t}^{*})-r_{t}(p)$ $\leq\tau\cdot a_{\max}p_{\max}+(2M+1)(1+N_{0}+N_{1})\cdot a_{\max}p_{\max}$ (16)
$\quad+\sum_{t=1}^{T}(\max\{C_{in},C_{out}\})(\frac{1}{N_{k_{t}}(t)}+\frac{1}{\tau})+\sum_{t=1}^{T}C_{non}\cdot\frac{1}{\tau}$
$(\text{ let }\tau=\sqrt{T})\rightarrow$ $=\tilde{O}(\sqrt{T}+T^{\frac{1}{4}}+\sum_{t=1}^{T}\frac{1}{N_{k_{t}}(t)}+\frac{T}{\tau})$
$=\tilde{O}(\sqrt{T}+\sum_{k=1}^{2M+1}\sum_{i_{k}=1}^{N_{k_{t}}(T)}\frac{1}{i_{k}})$
$=\tilde{O}(\sqrt{T}+T^{\frac{1}{4}}\log T)=\tilde{O}(\sqrt{T}).$
Here the first two rows are a decomposition of STAGE 1 (for $\tau$ rounds), STAGE 2 Initialization (for $2M+1$ rounds), STAGE 2 Case (a) and (b) (proposing $p_{t}=p_{k_{t}}$, Lemma˜4) and STAGE 2 Case (c) (proposing $p_{t}=\frac{\hat{a}}{2\hat{b}}$, Lemma˜5). The fourth row is by re-classification of $\frac{1}{N_{k_{t}}(t)}$ according to $k$, which leads to a summation over harmonic series (since each $N_{k_{t}}(t)$ increases by 1 for the same class $k_{t}=k$). By applying a union bound, Eq.˜16 holds with probability
@@
Here the first part comes from Lemma˜2, and the second part comes from Lemma˜3 for any $t\in[T]$ and $k\in\{-M,\ldots,M\}$. Let $\eta:=\frac{C_{a}+C_{b}\cdot p_{\max}}{10c\cdot T^{5/4}}$, and we show that Theorem˜5.1 holds. ∎
Remark 1
This $\tilde{O}(\sqrt{T})$ regret upper bound is near-optimal up to $\log{T}$ factors, as it matches the $\Omega(\sqrt{T})$ information-theoretic lower bound proposed by [10] for a no-censoring problem setting with linear noisy demand.
@@ 6 Discussions
Here we provide some insights on the current limitations and potential extension of this work to a broader field of research.
Extensions to Non-linear Demand Curve.
@@
In this work, we adopt a linear-and-noisy model for the potential demands, which is standard in dynamic pricing literature. Also, we utilize the unimodal property brought by this linear demand model, even after the censoring effect is imposed. If we generalize our methodologies to nonlinear demand functions, we have to carefully distinguish between potential local optima and saddle points that may also cause $r^{\prime}_{t}(p)=0$ for some sub-optimal $p$. We conjecture an $\Omega(T^{\frac{m+1}{2m+1}})$ lower bound in that case, where $m$ is the order of smoothness. It is worth investigating whether the censoring effect will introduce new local optimals or swipe off existing ones in multimodal settings.
Generalization to Unbounded Noises.
We assume the noise is bounded in a constant-width range. From the analysis in Section˜5, we know that the threshold of learning the optimal price in our problem setting is still the estimation of parameters. Therefore, this boundedness assumption streamlines the pure-exploration phase, facilitates the estimation of the parameters $b$ and $a$, and scales down the cumulative regret. While our methods and results can be extended to unbounded $O(\frac{1}{\log T})$-subGaussian noises by simple truncation, challenges remain for handling generic unbounded noises. Moreover, the problem can be more sophisticated with dual-censoring, both from above by inventory–as we have discussed– and from below by 0, especially when considering unbounded noises.
Extensions to Non-Lipschitz Noise CDF.
In this work, we assume the noise CDF as a Lipschitz function as many pricing-related works did [26, 52]. This assumption enables the local smoothness at $p_{t}^{*}$ and leads to a quadratic loss. However, this prevents us from applying our algorithm to non-Lipschitz settings, which even includes the noise-free setting. In fact, although we believe that a better regret rate exists for the noise-free setting, we have to state that the hardness of the problem is completely different with Lipschitz noises versus without it. Although a Lipschitz noise makes the observation “more blur”, it also makes the revenue curve “more smooth”. We would like to present an analog example from the feature-based dynamic pricing problem: When the Gaussian noise $\mathcal{N}(0,\sigma^{2})$ is either negligible (with $\sigma<\frac{1}{T}$, see [21]) or super significant (with $\sigma>1$, see [61]), the minimax regret is $O(\log T)$. However, existing works can only achieve $O(\sqrt{T})$ regret when $\sigma\in[\frac{1}{T},1]$. We look forward to future research on our problem setting once getting rid of the Lipschitzness assumption.
Extensions to Contextual Pricing.
In this work, we assume $a$ and $b$ are static, which may not hold in many real scenarios. Example˜1 serves as a good instance, showcasing significant fluctuations in popularity across different performances. A reasonable extension of our work would be modeling $a$ and $b$ as contextual parameters. Similar modelings have been adopt by [54] and [6] in the realm of personalized pricing research.
Societal Impacts.
@@
Our research primarily addresses a non-contextual pricing model that does not incorporate personal or group-specific data, thereby adhering to conventional fairness standards relating to temporal, group, demand and utility discrepancies as outlined by [20] and [17]. However, the non-stationarity of inventory levels could result in varying fulfillment rate over time, i.e. the proportions of satisfied demands at $\{p_{t}^{*}\}$’s might be different for $t=1,2,\ldots,T$. This raises concern regarding unfairness in fulfillment rate [48], particularly for products of significant social and individual importance.